The demand for tickets to an Ethiopian Camparada film is given by D(p)= 200,000-


10,000p, where p is the price of tickets.If the price of tickets is 12 birr, calculate price


elasticity of demand for tickets and draw the demand curve


1
Expert's answer
2022-01-25T17:13:26-0500

The price elasticity of demand measures the change in the quantity demanded of a good or service when the price of the product increases or decreases.


"\varepsilon_D=\dfrac{\%Change\ in \ demand}{\%Change\ in\ price\ unit}"

"=\dfrac{\Delta Q/\bar{Q}}{\Delta P/\bar{P}}"

"\dfrac{\Delta Q}{\Delta P}=-10000"

"D(12)=200000-10000(12)=80000"


"\varepsilon_D=-10000(\dfrac{12}{80000})=-1.5"

Therefore price elasticity of demand is -1.5.

As the price elasticity is negative it implies that as the price of the ticket increases demand will keep on falling.


"p=20-0.0001q"